# Leveraging LLMs, Graphs and Object Hierarchies for Task Planning in Large-Scale Environments

### v2

@@ Abstract
Planning methods struggle with computational intractability in solving task-level problems in large-scale environments. This work explores leveraging the commonsense knowledge encoded in LLMs to empower planning techniques to deal with these complex scenarios. We achieve this by efficiently using LLMs to prune irrelevant components from the planning problem’s state space, substantially simplifying its complexity. We demonstrate the efficacy of this system through extensive experiments within a household simulation environment, alongside real-world validation using a 7-DoF manipulator (video https://youtu.be/6ro2UOtOQS4).
I Introduction
In homes, hospitals, and factories, there is an increasing demand for robotic assistants that can react to high-level commands like “Clean room A.” Search-based planning [1, 2, 3] provides a promising framework for generating the action sequences necessary to solve such tasks. However, these techniques quickly reach their computational limits when applied to the realistic scenarios mentioned above, as these scenarios are populated by a large number of objects with which the robot can interact in multiple ways.
@@ I Introduction
To validate our method, experiments were conducted using the VirtualHome simulator [10], which contains environments with hundreds of objects, and a real-world 7-DoF robotic manipulator tasked with rearranging a configuration of objects on a table. This paper is structured as follows: Section II discusses related works in task planning and problem size reduction; Section III outlines the preliminary concepts and definitions; Section IV formulates the problem our method addresses; Section V details our proposed methodology; Section VI presents the experimental setup and results; and finally, Section VII concludes the paper and discusses future work directions.
II Related Work
@@
III Preliminaries
III-A Task Planning
@@
III-B Graph-based State Representation
@@ V Method
V-A Step 1: Environment-agnostic Object Selection
@@
V-A1 Taxonomy Graph States
@@
V-A2 Interacting with the taxonomy
@@
V-B Step 2: Relationship-based Object Selection
@@
V-B1 Iterative selection process
@@
V-C Integrating Taxonomy Graph States with LLMs
@@
For the function $LLM^{\mathcal{T}}$, it is only necessary to generate a string that lists the categories present at a given hierarchical level. For example, the following string could be included in the context window of the LLM: “Categories: electronics, lighting, appliances, …”. In the case of the lowest level, where object classes replace categories, the string could be: “Objects: apple, chair, toilet, …”.
For the relationship-based object step selection via $LLM^{\mathcal{R}}$, a similar structure is employed. Nevertheless, each component of the list consists of two objects and their corresponding relationship. For instance, “Relationships between objects: banana (34) is inside the fridge (70), plate (67) is on the table (45), …”. Here, the numbers in parentheses correspond to the unique IDs of individual objects. These are necessary because relationships pertain to specific objects rather than object classes. In this example, the distinction ensures that not every banana is assumed to be inside every fridge, but rather that only the banana with ID 34 is inside the fridge with ID 70.
V-D Planning with State Graphs
@@
V-D1 Search-based Policies
@@ VI Experiments
VI-A State-space size reduction
TABLE I: Number of selected objects with GPT-4o and GPT-3.5 on successful trials. ‘N’ denotes the number of subtasks.
@@ VI-A State-space size reduction
VI-A1 Unsuccessful state-space reductions
@@
VI-A2 Size of successful reductions
Apart from evaluating whether the selected objects contain $O_{\mathrm{g}}$, it is also important to study the size of the set of selected objects $\bar{O}$. To this end, Table I provides the average number of selected objects for the successful cases of Fig. 4. This table also shows the average initial number of objects and the average number of necessary objects. For both GPT-4o and GPT-3.5, we can observe that $|\bar{O}|$ is similar to $|O_{\mathrm{g}}|$, especially when compared to the initial number of objects, which is orders of magnitude larger. It is interesting to note that in multiple cases, GPT-3.5 selects fewer objects than GPT-4o. At first glance, this might indicate better performance for GPT-3.5. However, this is because GPT-3.5 fails to properly extract some information from the task goal and graph during the object selection process, leading to more simplistic solutions that overlook relevant objects and, consequently, to the high failure rates observed in Fig. 4.
TABLE II: Comparison of success rates and average trajectory steps across different methods using the obtained reduced states.
@@ VI-A2 Size of successful reductions
VI-B Planning
Given that the objective of reducing the state space size is to simplify the planning problem, in Table II, we provide the performance of LLM-based planners and a search-based approach, namely, Monte Carlo Tree Search (MCTS) [17], in reduced-state-space settings. The employed LLM-based planners operate on a per-time-step basis, as explained in Section V-D. Every planner was executed over the reduced states obtained using GPT-4o.
VI-B1 Planning performance with state space reduction
From Table II, we observe that GPT-4o outperforms GPT-3.5 and MCTS in planning, achieving a 0.73 success rate even in the most challenging scenarios. The performance of GPT-3.5 is surprisingly low; we noted that this occurred because the model could not properly reason from the provided context, thereby selecting actions that did not bring the agent closer to the goal. Lastly, MCTS obtained the expected performance: as the problem size increases, it becomes increasingly intractable. It is also worth noting that the number of steps in the solutions found by MCTS is considerably larger than those found by the LLM-based planners. This result is expected, as the primary objective of MCTS is to find a solution, not necessarily the best solution, which can lead to redundancies or unnecessary steps. In contrast, LLM-based planners provided solutions guided by their commonsense knowledge.
VI-B2 Planning performance without state space reduction
The same experiments shown in Table II were conducted for MCTS without reducing the state space size. As expected, MCTS was unable to find any solutions in these scenarios, not even for problems with only one subtask, highlighting the importance of reducing the state space size. For LLM-based planners, obtaining such results was not feasible, as planning with such long context windows would be prohibitively expensive. This further demonstrates the advantage of reducing the state space size.
VI-B3 Overall performance
@@
VI-C Real-world validation

### v3

# Scalable Task Planning via Large Language Models and Structured World Representations

@@ Abstract
Abstract— Planning methods struggle with computational intractability in solving task-level problems in large-scale environments. This work explores leveraging the commonsense knowledge encoded in LLMs to empower planning techniques to deal with these complex scenarios. We achieve this by efficiently using LLMs to prune irrelevant components from the planning problem’s state space, substantially simplifying its complexity. We demonstrate the efficacy of this system through extensive experiments within a household simulation environment, alongside real-world validation using a 7-DoF manipulator (video https://youtu.be/6ro2UOtOQS4).
keywords:
Task Planning, LLMs, Graphs, Taxonomy, Robotics
Rodrigo Pérez-Dattari1, Zhaoting Li1, Robert Babuška1,2 Jens Kober1 and Cosimo Della Santina1,3,
{affiliations}
1Department of Cognitive Robotics, Delft University of Technology, Netherlands.
2Czech Institute of Informatics, Robotics, and Cybernetics, Czech Technical University in Prague, Czech Republic.
3Institute of Robotics and Mechatronics, German Aerospace Center (DLR), Wessling, Germany.
1 Introduction
In homes, hospitals, and factories, there is an increasing demand for robotic assistants that can react to high-level commands like “Clean room A." Search-based planning [1, 2, 3] provides a promising framework for generating the action sequences necessary to solve such tasks. However, these techniques quickly reach their computational limits when applied to the realistic scenarios mentioned above, as these scenarios are populated by a large number of objects with which the robot can interact in multiple ways.
@@ 1 Introduction
To validate our method, experiments were conducted using the VirtualHome simulator [10], which contains environments with hundreds of objects, and a real-world 7-DoF robotic manipulator tasked with rearranging a configuration of objects on a table. This paper is structured as follows: Section 2 discusses related works in task planning and problem size reduction; Section 3 outlines the preliminary concepts and definitions; Section 4 formulates the problem our method addresses; Section 5 details our proposed methodology; Section 6 presents the experimental setup and results; and finally, Section 7 concludes the paper and discusses future work directions.
2 Related Work
@@
3 Preliminaries
3.1 Task Planning
@@
3.2 Graph-based State Representation
@@ 5 Method
5.1 Step 1: Environment-agnostic Object Selection
@@
5.1.1 Taxonomy Graph States
@@
5.1.2 Interacting with the taxonomy
@@
5.2 Step 2: Relationship-based Object Selection
@@
5.2.1 Iterative selection process
@@
5.3 Integrating Taxonomy Graph States with LLMs
@@
For the function $LLM^{\mathcal{T}}$, it is only necessary to generate a string that lists the categories present at a given hierarchical level. For example, the following string could be included in the context window of the LLM: “Categories: electronics, lighting, appliances, …". In the case of the lowest level, where object classes replace categories, the string could be: “Objects: apple, chair, toilet, …".
For the relationship-based object step selection via $LLM^{\mathcal{R}}$, a similar structure is employed. Nevertheless, each component of the list consists of two objects and their corresponding relationship. For instance, “Relationships between objects: banana (34) is inside the fridge (70), plate (67) is on the table (45), …". Here, the numbers in parentheses correspond to the unique IDs of individual objects. These are necessary because relationships pertain to specific objects rather than object classes. In this example, the distinction ensures that not every banana is assumed to be inside every fridge, but rather that only the banana with ID 34 is inside the fridge with ID 70.
5.4 Planning with State Graphs
@@
5.4.1 Search-based Policies
@@ 6 Experiments
Table 1: Number of selected objects with GPT-4o and GPT-3.5 on successful trials. ‘N’ denotes the number of subtasks.
@@ 6.1 State-space size reduction
6.1.1 Unsuccessful state-space reductions
@@
6.1.2 Size of successful reductions
Apart from evaluating whether the selected objects contain $O_{\mathrm{g}}$, it is also important to study the size of the set of selected objects $\bar{O}$. To this end, Table 1 provides the average number of selected objects for the successful cases of Fig. 4. This table also shows the average initial number of objects and the average number of necessary objects. For both GPT-4o and GPT-3.5, we can observe that $|\bar{O}|$ is similar to $|O_{\mathrm{g}}|$, especially when compared to the initial number of objects, which is orders of magnitude larger. It is interesting to note that in multiple cases, GPT-3.5 selects fewer objects than GPT-4o. At first glance, this might indicate better performance for GPT-3.5. However, this is because GPT-3.5 fails to properly extract some information from the task goal and graph during the object selection process, leading to more simplistic solutions that overlook relevant objects and, consequently, to the high failure rates observed in Fig. 4.
Table 2: Comparison of success rates and average trajectory steps across different methods using the obtained reduced states.
@@ 6.1.2 Size of successful reductions
6.2 Planning
Given that the objective of reducing the state space size is to simplify the planning problem, in Table 2, we provide the performance of LLM-based planners and a search-based approach, namely, Monte Carlo Tree Search (MCTS) [17], in reduced-state-space settings. The employed LLM-based planners operate on a per-time-step basis, as explained in Section 5.4. Every planner was executed over the reduced states obtained using GPT-4o.
6.2.1 Planning performance with state space reduction
From Table 2, we observe that GPT-4o outperforms GPT-3.5 and MCTS in planning, achieving a 0.73 success rate even in the most challenging scenarios. The performance of GPT-3.5 is surprisingly low; we noted that this occurred because the model could not properly reason from the provided context, thereby selecting actions that did not bring the agent closer to the goal. Lastly, MCTS obtained the expected performance: as the problem size increases, it becomes increasingly intractable. It is also worth noting that the number of steps in the solutions found by MCTS is considerably larger than those found by the LLM-based planners. This result is expected, as the primary objective of MCTS is to find a solution, not necessarily the best solution, which can lead to redundancies or unnecessary steps. In contrast, LLM-based planners provided solutions guided by their commonsense knowledge.
6.2.2 Planning performance without state space reduction
The same experiments shown in Table 2 were conducted for MCTS without reducing the state space size. As expected, MCTS was unable to find any solutions in these scenarios, not even for problems with only one subtask, highlighting the importance of reducing the state space size. For LLM-based planners, obtaining such results was not feasible, as planning with such long context windows would be prohibitively expensive. This further demonstrates the advantage of reducing the state space size.
6.2.3 Overall performance
@@
6.3 Real-world validation